Sue now has $125. How much would she have after 8 years if she leaves it...

  1. Sue now has $125. How much would she have after 8 years if she leaves it invested at 8.5% with annual compounding?
  1. $205.83
  2. $216.67
  3. $228.07
  4. $240.08
  5. $252.08
  1. Suppose the U.S. Treasury offers to sell you a bond for $3,000. No payments will be made until the bond matures 10 years from now, at which time it will be redeemed for $5,000. What interest rate would you earn if you bought this bond at the offer price?
  1. 3.82%
  2. 4.25%
  3. 4.72%
  4. 5.24%
  5. 5.77%
  1. What’s the present value of a perpetuity that pays $250 per year if the appropriate interest rate is 5%?
  1. $4,750
  2. $5,000
  3. $5,250
  4. $5,513
  5. $5,788
  1. What’s the rate of return you would earn if you paid $950 for a perpetuity that pays $85 per year?
  1. 8.95%
  2. 9.39%
  3. 9.86%
  4. 10.36%
  5. 10.88%
Answer #1

1) future value = present value(1+r)^n

where , r = rate of interest

n = number of years

so future value = 125(1.085)^8

= 125 x 1.9206043

= $240.08

2)

using the above formula we have to find ' r '

we know present value = 3000

future value = 5000

5000 = 3000(1 + r)^10

(1+r)^10 = 1.66667

r = (1.66667)^1/10 - 1

r = 0.0524 or 5.24%

3)

present value = yearly payment / interest rate

present value = 250 / 5%

= $5000

4)

using the above formula we have to find out interest rate

950 = 85 / r

so r = 85 / 950

= 0.08947 or 8.95%